the length of a rectangular garden is 3 yd more than twice its width. the perimeter of the garden is 36 yd. what are the width and length of the garden?

Answer by Timnewman(323)

Hi dear,
L=lenght
w=width
From your question,
L=2w+3----(1)
2(L+w)=36
L+w=18-----(2)
put 1 in 2
2w+3+w=18
3w=18-3
3w=15
w=5
put w=5 in equ 1
L=2(5)+3
L=10+3
L=13
The lenght is 13yd,
Width=5yd.
Cheers;
Tim.